# Graph Neural Networks Gone Hogwild

### v1

@@ 6.3 Results
Table 2: Decrease in task performance (decrease in accuracy for CHAINS, MNIST, and increase in relative RMSE for COUNT, SUM, COORDINATES) observed from switching from synchronous to asynchronous inference on sub-sample of test data (10 samples) using one trained model instance. Mean and standard deviation are across 5 asynchronous runs. The poor performance of GCN and GAT are consistent with the expected unreliability of explicitly-defined GNNs with asynchronous inference.
@@
IGNN $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
GSDGNN $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Energy GNN node-wise $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Energy GNN edge-wise $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Energy GNN + attention $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
@@
In Table 2, we demonstrate empirically that the explicitly-defined architectures GCN and GAT perform poorly and unreliably under asynchrony, with task performance decreasing for all experiments. In our experiments, we simulate asynchronous inference; our algorithm is in Appendix E. In most cases the variance of the performance decrease is large as a result of inconsistent predictions under different random node update schedules. In cases where the variance is low, we observe that predictions for different random schedules collapse to the same/similar values due to non-linearities in the architecture. We reiterate that explicitly-defined GNNs are not suitable for the tasks we consider, in which we assume asynchronous, distributed inference and aim for reliable predictions under this regime. For implicitly-defined GNNs, the decrease in performance under asynchronous inference is less than 0.1%, empirically confirming the convergence guarantees given by Proposition 4.1 and Proposition 4.2.
@@ 7 Related work
Asynchronous algorithms
Asynchronous algorithmic models (sometimes called chaotic relaxation models) date back at least to the 1960s (Chazan and Miranker, 1969), and were explored extensively into the 1970s and 1980s (Donnelly, 1971; Miellou, 1975; Robert et al., 1975; Baudet, 1978; Bertsekas, 1982, 1983; Bojańczyk, 1984; Mitra, 1987; Üresin and Dubois, 1989). In Section 4 we study GNNs under partial asynchrony, a model which imposes constraints on the sequencing of computations and frequency of communication between distributed elements. Our analysis of partially asynchronous GNN inference draws directly from prior work which analyzes the sufficient conditions for convergence (Tsitsiklis, 1984; Tsitsiklis et al., 1986; Bertsekas and Tsitsiklis, 1989).
Asynchronous GNN training
There has been a plethora of work in distributed training of GNNs, where data are computed on by separate workers (Besta and Hoefler, 2022; Shao et al., 2022). In this regime, workers may not be operating on independent data, e.g., when data come from a single connected graph. Since ignoring this dependence reduces performance, workers exchange embedding information associated with “boundary vertices” which are logically connected but delegated to different workers. Some distributed training assumes workers operate asynchronously, using stale embedding (Md et al., 2021; Peng et al., 2022; Wan et al., 2022) or gradient information (Thorpe et al., 2021) corresponding to previous training epochs. However, across all these frameworks, the forward pass is executed synchronously, and to our knowledge there is no work examining asynchronous GNN inference.
GNNs applied to multi-agent tasks
Several works have identified GNNs as a useful tool in parameterizing decentralized policies for multi-agent systems, with applications including flock formation, target tracking, path planning, and channel allocation in wireless networks. These works employ GNNs to generate neighborhood-aggregated state representations for agent control policies, which are optimized through imitation learning (Grattarola et al., 2021; Zhou et al., 2021; Gama et al., 2020) or multi-agent reinforcement learning (Nakashima et al., 2019; Khan et al., 2019; Jiang and Lu, 2018; Jiang et al., 2020). A pitfall of these works is that they do not consider asynchronous execution of the GNNs which generate state representations. Consistent with the observations in our work, Blumenkamp et al. (2021) show that performance of multi-agent policies that use GNNs suffer in real-world deployment where synchronous execution is not possible.
@@ 8 Conclusion
GNNs have the potential to provide learning frameworks for distributed systems, with applications to privacy, robotics, remote sensing, and other domains. However, most conventional GNN architectures are not compatible with asynchronous inference and this hinders their deployment on these applications.
In this work, we characterize the class of implicitly-defined GNNs as being provably robust to partially asynchronous inference. Motivated by the lack of diversity in this class of architectures, we contribute a novel addition in the form of energy GNNs. We show that energy GNNs achieve better performance than other implicitly-defined GNNs on a number of synthetic tasks motivated by decentralized and constrained multi-agent systems, where distributed, asynchronous inference is desirable. In addition to its robustness to asynchrony, our method is comparable in generalization performance (on benchmarks) with other modern GNN architectures that do not offer these guarantees.
The positive results of our synthetic experiments motivates additional work in applying GNN architectures to multi-agent related tasks. A specific line of work which we expect to be interesting is real-time inference of dynamic graphs, both because of the relevance to problems in, e.g., robotics, but also due to the “anytime” nature of implicitly-defined GNN architectures. Distributed and asynchronous inference over large graphs is another application for implicitly-defined GNNs which should be explored. We are optimistic that the energy GNN framework itself, and implicitly-defined GNNs in general, will provide a path forward for “learning to learn” distributed, asynchronous algorithms.
There are several limitations of this work. The forward pass in implicitly-defined GNNs is more involved computationally, and may require an amount of time that cannot be predetermined. The condition number of the PICNN has a significant effect on convergence rate and is difficult to control. The backward pass requires implicit gradients rather than simple backpropagation. Our model for partial asynchrony, although standard, makes assumptions that may not match real-world systems. Finally, operationalizing the theoretical results may require constants that are not readily available.
@@ Appendix A GNN architectures
GCNs (Kipf and Welling, 2017) replace the adjacency matrix ${\bm{A}}$ with the symmetric normalized adjacency matrix with added self-loops, $\tilde{{\bm{A}}}=({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}({\bm{A}}+{\bm{I}})({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}$. With node embeddings initialized to be equal to the node features, $f_{\theta}^{\ell}$ is defined as:
@@
GATs (Veličković et al., 2018) apply an attention mechanism to determine the weighting of information from different neighbors. With node embeddings initialized to be equal to the node features, $f_{\theta}^{\ell}$ is defined as:
@@
IGNNs Gu et al. (2020) are a fixed point GNN architecture, in which each parameterized node update layer is contractive with respect to the node embeddings. For simplicity, we assume a single parameterized layer $f_{\theta}$ and exclude the layer superscript from our notation. The parameterized node embedding update is repeated for steps $t=0,...,T-1$, where the stopping point $T$ is determined by when node embeddings converge (within some numerical tolerance). IGNNs use a similar embedding update function as GCN, but add node features as an additional input to the update function. A layer $f_{\theta}$ is defined as:
${\bm{m}}_{i}(t)$ $:=\textstyle\sum_{j\in\mathcal{N}(i)}\tilde{{\bm{A}}}_{i,j}\theta_{m}{\bm{h}}_{j}(t)$ ${\bm{h}}_{i}(t+1)$ $:=\phi({\bm{m}}_{i}(t)+g({\bm{x}}_{i};\theta_{u})),$ (21)
where ${\theta_{m}\in\mathbb{R}^{k\times k}}$,  $g_{\theta}:\mathbb{R}^{n\times p}\rightarrow\mathbb{R}^{n\times k}$ and $\phi$ is a component-wise non-expansive function such as ReLU. Convergence is guaranteed by constraining $||\theta_{m}||_{\infty}<\lambda_{max}(\tilde{{\bm{A}}})^{-1}$, where $\lambda_{max}$ is the maximum eigenvalue of $\tilde{{\bm{A}}}$. This ensures that the update is contractive, a sufficient condition for convergence. Since the fixed point is unique, ${\bm{h}}_{i}(0)$ can be initialized arbitrarily (although convergence time will vary).
@@
EIGNNs Liu et al. (2021) are another fixed point GNN architecture which are very similar to IGNNs. The message passing iteration is constructed such that a closed-form solution can be obtained for the fixed point of a layer, which is more efficient than iterating message passing to convergence. A layer $f_{\theta}$ is defined as:
@@
Scarselli et al. (2009) introduce a general nonlinear fixed point GNN whose update can be written as
@@ Appendix B Input-convex GNN architecture details
As in Amos et al. (2017), we construct a parametric family of neural networks $f_{\theta}({\bm{x}},{\bm{y}})$ with inputs ${\bm{x}}\in\mathbb{R}^{n},{\bm{y}}\in\mathbb{R}^{m}$ which are convex with respect to ${\bm{y}}$ (i.e. a subset of the inputs). We define a $k$-layer partially convex neural network by the recurrences:
${\bm{u}}_{i+1}$ $=\tilde{g}_{i}(\tilde{\bm{W}}_{i}{\bm{u}}_{i}+\tilde{\bm{b}}_{i})$
${\bm{z}}_{i+1}$ $=g_{i}($
${\bm{W}}_{i}^{(z)}({\bm{z}}_{i}\circ[{\bm{W}}_{i}^{(zu)}{\bm{u}}_{i}+{\bm{b}}_{i}^{(z)}]_{+})+$
${\bm{W}}_{i}^{(y)}({\bm{y}}\circ({\bm{W}}_{i}^{(yu)}{\bm{u}}_{i}+{\bm{b}}_{i}^{(y)}))+$
${\bm{W}}_{i}^{(u)}{\bm{u}}_{i}+{\bm{b}}_{i})$
$f_{\theta}({\bm{x}},{\bm{y}})={\bm{z}}_{k},\quad{\bm{u}}_{0}={\bm{x}},\quad{\bm{z}}_{0}=\mathbf{0}$
Provided the $W^{(z)}$ are elementwise nonnegative, and the activation functions $g_{i}$ are non-decreasing in each argument, it follows that $f_{\theta}$ is convex in ${\bm{y}}$.
In the context of an energy GNN, $E_{\theta}$ is comprised of two PICNNs, specialized to operate on graph structures; we refer to the resulting architecture as a partially input-convex GNN (PICGNN). The message function $m$ in Equation 16 corresponds to a PICNN which has node embeddings ${\bm{h}}_{i},i=1,\dots,n$ as its convex inputs (all other features are non-convex inputs). A second PICNN corresponds with the update function $u$ in Equation 15, which is convex in the node embeddings and the messages computed by the message function, and nonconvex in the node features. The graph energy $E_{\theta}$ can then be written as the sum of the outputs of the second updating PICNN.
@@ Appendix D Proof of Convergence Under Partial Asynchrony
Our guarantee of convergence uses the classical result from Bertsekas and Tsitsiklis (1989, Chapter 7.5), which depends on several formal assumptions. We reproduce these assumptions here, in the notation used in the present paper, and address how they are satisfied in our problem setting. In the following, $s_{i}(t)$ is the “search direction” taken by node $i$, i.e., the vector used by node $i$ to take optimization step. Ideally, this would be the negative gradient $-\nabla_{{\bm{h}}_{i}}E_{\theta}$, but partial asynchrony means it may be a vector constructed from stale information. Our goal is to show that convergence of the optimization is nevertheless guaranteed. We use ${\bm{h}}\in\mathbb{R}^{nk}$ to denote the unrolled embeddings ${\bm{H}}$.
Assumption D.1 (Bertsekas and Tsitsiklis (1989) Assumption 5.1).
@@
Assumption D.2 (Bertsekas and Tsitsiklis (1989) Assumption 5.5).
@@ Appendix E Asynchronous GNN implementation
In our asynchronous inference experiments, we simulate partially asynchronous execution (see Algorithm 1). We fix the maximum staleness bound to $B=7$. Staleness is achieved by staggering node updates across time and introducing delay in messages sent. In particular, when a node updates, its next update time is selected randomly between time $t+1$ and $t+S$, where $S$ is the “stagger” time, and one of the last $D$ values of its neighbors is chosen for performing the node update. This satisfies assumptions 1 and 2 of partial asynchronism; nodes update at least every $S$ time units, and a node’s view of its neighbors is stale by at most $B=S+D$ time steps.
@@
Initialize each node in ${\mathcal{G}}$ with all GNN parameters, its own node features ${\bm{X}}_{i}$, and for each of its neighbors $j$, the node and edge features ${\bm{X}}_{j}$, ${\bm{E}}_{ij}$, and weights ${\bm{A}}_{i,j}$. Let $L$ be the number of layers in the GNN, equal to $\infty$ for implicitly-defined GNNs. Let $T$ be the total number of simulated node updates. Let $n$ be the number of nodes in the graph. Let $B=S+D$ be the maximum staleness of a node’s view of its neighbor values.
@@
${\bm{m}}_{i}\leftarrow\text{\small AGGREGATE}^{t_{i}}\left(\{m^{t_{i}}({\bm{h}}_{j}(\tau_{j}^{i}),{\bm{h}}_{i},{\bm{x}}_{j},{\bm{x}}_{i},{\bm{e}}_{ij};\theta_{m}^{t_{i}})\mid j\in\mathcal{N}(i)\},{\bm{A}}_{i};\theta_{g}^{t_{i}}\right)$

### v2

@@ 6.3 Results
Table 2: Decrease in task performance (decrease in accuracy for CHAINS, MNIST, and increase in relative RMSE for COUNT, SUM, COORDINATES) observed from switching from synchronous to asynchronous inference on sub-sample of test data (10 samples) using one trained model instance. Mean and standard deviation are across 5 asynchronous runs. The poor performance of GCN and GAT are consistent with the expected unreliability of explicitly-defined GNNs with asynchronous inference. Decreases in task performance for all implicitly-defined GNNs (IGNN, GSDGNN, and energy GNN variants) is less than $0.1\%$ (i.e. result from numerical error); these values are omitted from the table for concision.
@@
In Table 2, we demonstrate empirically that explicitly-defined architectures such as GCN and GAT perform poorly and unreliably under asynchrony, with task performance decreasing for all experiments. In our experiments, we simulate asynchronous inference; our algorithm is in Appendix E. In most cases the variance of the performance decrease is large as a result of inconsistent predictions under different random node update schedules and communication delays. In cases where the variance is low, we observe that predictions for different random schedules collapse to the same/similar values due to non-linearities in the architecture. For implicitly-defined GNNs, the decrease in performance under asynchronous inference is less than 0.1%, empirically confirming the convergence guarantees given by Proposition 4.1 and Proposition 4.2.
@@ 7 Related work
Dudzik et al. (2023) axiomatically derive GNN layers which are invariant to asynchrony. For instance, using max-aggregation for messages is by construction invariant to asynchrony; messages from neighbors can be processed as they arrive rather than waiting for all inputs to be available, and the final aggregated message will still be correct. Later work in this thread aims at enforcing asyncrony invariance through a self-supervised loss function rather than using layers which are asynchrony invariant by construction (Monteagudo-Lago et al., 2024). The motivation for this line of work is to apply GNNs for neural algorithmic reasoning, where the GNN is trained to emulate classical algorithms such as the Bellman-Ford algorithm. Enforcing asyncrony invariance in the GNN serves to improve alignment between the algorithm to be learned and the GNN computation. The asynchronous model of computation used in these works differs from ours in that nodes must maintain information from all layers in the GNN, rather than executing layer-wise.
Faber and Wattenhofer (2024) also explore asynchronous communication between nodes during GNN inference. Their approach consists of processing node updates in a queue, rather than applying all node updates in a layer in parallel. The goal of their work is to leverage sequential node updates to improve under-reaching and over-squashing in learned node embeddings. Asynchronous execution in their model is deterministic in that there are no random delays in communication between nodes; nodes receive messages and perform updates in the same order across different runs. They show that random delays and node update orderings leads to unstable training and results in decreased task performance. This makes their approach unsuitable for the settings we consider, in which asynchrony is not controlled.
@@ 8 Conclusion
GNNs have the potential to provide learning frameworks for decentralized multi-agent systems, with applications to robotics, remote sensing, and other domains. However, asynchronous execution and unreliable communication between agents are common features in real-world deployment, and conventional GNN architectures do not generate reliable predictions in this setting. In this work, we characterize the class of implicitly-defined GNNs as being provably robust to partially asynchronous inference. Motivated by lack of diversity in this class of architectures, we contribute a novel addition in the form of energy GNNs, which achieve better performance than other implicitly-defined GNNs on a number of synthetic multi-agent tasks.
The positive results of our synthetic experiments motivates additional work in applying GNN architectures to multi-agent systems, particularly making use of implicitly-defined GNNs to generate state representations for control tasks, rather than for prediction tasks. A specific line of work which we expect to be interesting is real-time inference on dynamic graphs because of the relevance to problems in, e.g., robotics. Distributed and asynchronous inference for datasets consisting of large graphs, where inference must be distributed among multiple processors due to scale, is another application for implicitly-defined GNNs which should be explored.
There are several limitations of this work, particularly in training implicitly-defined GNNs. The forward pass requires convergence of node embeddings via a fixed point iteration, and requires an amount of time/computation that cannot be known in advance. During training, we decrease the number of iterations required for convergence by initializing node embeddings to those from the previous epoch of training. Real-time inference in multi-agent systems is likely to also enjoy the benefits of a warm-start initialization. This is because the graph features and structure are often expected to change slowly over time, meaning that after an initial “boot-up” from random initialization, a continuously operating implicitly-defined GNN is likely to have embeddings initialized close to the solution from one time step to the next (assuming the node embeddings ${\bm{H}}*$ are not extremely sensitive to small changes in the graph). For optimization-based GNNs using gradient-based optimization in the forward pass, the condition number of the Hessian of $E_{\theta}$ affects the convergence rate, but is difficult to control. Similarly, if $F_{\theta}$ is poorly conditioned in fixed-point GNNs, convergence can be slow and is susceptible to numerical instability. During training, we use implicit differentiation to obtain gradients w.r.t. parameters rather than simple backpropagation; this avoids unrolling through the iterative process in the forward pass, but requires solving a linear system using the Hessian of $E_{\theta}$ (for optimization-based GNNs) or the Jacobian of $F_{\theta}$ (for fixed-point GNNs), either of which may be poorly conditioned. Finally, operationalizing the theoretical results may require constants that are not readily available.
@@ Appendix A GNN architectures
GCNs [Kipf and Welling, 2017] replace the adjacency matrix ${\bm{A}}$ with the symmetric normalized adjacency matrix with added self-loops, $\tilde{{\bm{A}}}=({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}({\bm{A}}+{\bm{I}})({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}$. With node embeddings initialized to be equal to the node features, $f_{\theta}^{\ell}$ is defined as:
@@
GATs [Veličković et al., 2018] apply an attention mechanism to determine the weighting of information from different neighbors. With node embeddings initialized to be equal to the node features, $f_{\theta}^{\ell}$ is defined as:
@@
IGNNs Gu et al. [2020] are a fixed point GNN architecture, in which each parameterized node update layer is contractive with respect to the node embeddings. For simplicity, we assume a single parameterized layer $f_{\theta}$ and exclude the layer superscript from our notation. The parameterized node embedding update is repeated for steps $t=0,...,T-1$, where the stopping point $T$ is determined by when node embeddings converge (within some numerical tolerance). IGNNs use a similar embedding update function as GCN, but add node features as an additional input to the update function. A layer $f_{\theta}$ is defined as:
${\bm{m}}_{i}(t)$ $:=\textstyle\sum_{j\in\mathcal{N}(i)}\tilde{{\bm{A}}}_{i,j}\theta_{m}{\bm{h}}_{j}(t)$ ${\bm{h}}_{i}(t+1)$ $:=u({\bm{m}}_{i}(t)+g({\bm{x}}_{i};\theta_{u})),$ (22)
where ${\theta_{m}\in\mathbb{R}^{k\times k}}$,  $g_{\theta}:\mathbb{R}^{n\times p}\rightarrow\mathbb{R}^{n\times k}$ and $\phi$ is a component-wise non-expansive function such as ReLU. Convergence is guaranteed by constraining $||\theta_{m}||_{\infty}<\lambda_{pf}(\tilde{{\bm{A}}})^{-1}$, where $\lambda_{pf}(\tilde{A})$ is the maximum eigenvalue of $\tilde{|{\bm{A}}|}$. This ensures that the update is contractive, a sufficient condition for convergence. Since the fixed point is unique, ${\bm{h}}_{i}(0)$ can be initialized arbitrarily (although convergence time will vary).
@@
EIGNNs Liu et al. [2021] are another fixed point GNN architecture which are very similar to IGNNs. The message passing iteration is constructed such that a closed-form solution can be obtained for the fixed point of a layer, which is more efficient than iterating message passing to convergence. A layer $f_{\theta}$ is defined as:
@@
Scarselli et al. [2009] introduce a general nonlinear fixed point GNN whose update can be written as
@@ Appendix B Input-convex GNN architecture details
As in Amos et al. [2017], we construct a parametric family of neural networks $f_{\theta}({\bm{x}},{\bm{y}})$ with inputs ${\bm{x}}\in\mathbb{R}^{n},{\bm{y}}\in\mathbb{R}^{m}$ which are convex with respect to ${\bm{y}}$ (i.e. a subset of the inputs). An $L$-layer partially convex neural network is defined by the following recurrences:
${\bm{u}}_{\ell+1}$ $=\tilde{g}_{\ell}(\tilde{\bm{W}}_{\ell}{\bm{u}}_{\ell}+\tilde{\bm{b}}_{\ell})$
${\bm{z}}_{\ell+1}$ $=g_{\ell}($
${\bm{W}}_{\ell}^{(z)}({\bm{z}}_{\ell}\circ[{\bm{W}}_{\ell}^{(zu)}{\bm{u}}_{\ell}+{\bm{b}}_{\ell}^{(z)}]_{+})+$
${\bm{W}}_{\ell}^{(y)}({\bm{y}}\circ({\bm{W}}_{\ell}^{(yu)}{\bm{u}}_{\ell}+{\bm{b}}_{\ell}^{(y)}))+$
${\bm{W}}_{\ell}^{(u)}{\bm{u}}_{\ell}+{\bm{b}}_{\ell})$
$f_{\theta}({\bm{x}},{\bm{y}})={\bm{z}}_{L},\quad{\bm{u}}_{0}={\bm{x}},\quad{\bm{z}}_{0}=\mathbf{0}$
Provided the ${\bm{W}}^{(z)}$ are elementwise nonnegative for all layers $\ell$, and the activation functions $g_{\ell}$ are non-decreasing in each argument, it follows that $f_{\theta}$ is convex in ${\bm{y}}$.
In the context of an energy GNN, $E_{\theta}$ is comprised of two PICNNs, specialized to operate on graph structures; we refer to the resulting architecture as a partially input-convex GNN (PICGNN). The message function $m$ in Equation 17 corresponds to a PICNN which has node embeddings ${\bm{h}}_{i},i=1,\dots,n$ as its convex inputs (all other features are non-convex inputs). A second PICNN corresponds with the update function $u$ in Equation 16, which is convex in the node embeddings and the messages computed by the message function, and nonconvex in the node features. The graph energy $E_{\theta}$ can then be written as the sum of the outputs of the second PICNN.
@@ Appendix D Proof of Convergence Under Partial Asynchrony
Our guarantee of convergence uses the classical result from Bertsekas and Tsitsiklis [1989, Chapter 7.5], which depends on several formal assumptions. We reproduce these assumptions here, in the notation used in the present paper, and address how they are satisfied in our problem setting. In the following, $s_{i}(t)$ is the “search direction” taken by node $i$, i.e., the vector used by node $i$ to take optimization step. Ideally, this would be the negative gradient $-\nabla_{{\bm{h}}_{i}}E_{\theta}$, but partial asynchrony means it may be a vector constructed from stale information. Our goal is to show that convergence of the optimization is nevertheless guaranteed. We use ${\bm{h}}\in\mathbb{R}^{nk}$ to denote the unrolled embeddings ${\bm{H}}$.
Assumption D.1 (Bertsekas and Tsitsiklis [1989] Assumption 5.1).
@@
Assumption D.2 (Bertsekas and Tsitsiklis [1989] Assumption 5.5).
@@ Appendix E Asynchronous GNN implementation
In our asynchronous inference experiments, we simulate partially asynchronous execution (see Algorithm 1). We fix the maximum staleness bound to $B=5$. Node updates are staggered across time and messages sent between nodes can incur delays. In particular, when a node updates, its next update time is selected randomly between time $t+1$ and $t+S$, where $S:=5$ is the “stagger” time, and one of the last $D:=2$ values of its neighbors is chosen for performing the node update. This satisfies assumptions 1 and 2 of partial asynchronism; nodes update at least every $S$ time units, and messages a node’s view of its neighbors is stale due to message delay by at most $D$ time steps.
@@
Initialize each node in ${\mathcal{G}}$ with all GNN parameters, its own node features ${\bm{X}}_{i}$, and for each of its neighbors $j$, the node and edge features ${\bm{X}}_{j}$, ${\bm{E}}_{ij}$, and weights ${\bm{A}}_{i,j}$. Let $L$ be the number of layers in the GNN, equal to $\infty$ for implicitly-defined GNNs. Let $T$ be the total number of simulated node updates. Let $n$ be the number of nodes in the graph. Let $S$ be the maximum number of time units between updates for any node, and $D$ be the maximum delay incurred by messages from neighbors.
@@
${\bm{m}}_{i}\leftarrow\bigoplus_{j\in\mathcal{N}(i)}m^{t_{i}}\left({\bm{h}}_{i},{\bm{h}}_{j}(\tau_{j}^{i}(t)),{\bm{x}}_{i},{\bm{x}}_{j},{\bm{e}}_{ij};\theta_{m}^{t_{i}}\right)$